# Behind the Myth of Exploration in Policy Gradients

### v2

@@ 5 Conclusion
In conclusion, this research takes a step towards dispelling misunderstandings about exploration through the study of its effects on the performance of policy-gradient algorithms. More particularly, we distinguished two effects exploration has on the optimization. First, it modifies the learning objective in order to remove local extrema. Second, it modifies the gradient estimates and increases the likelihood that the update steps lead to improved returns. These two phenomena were studied through four criteria that we introduced and illustrated.
@@
These ideas apply to other direct policy optimization algorithms. Indeed, the four criteria do not assume any structure on the learning objective and can thus be straightforwardly applied to any objective function optimized by a direct policy search algorithm. In particular, for off-policy policy gradient, we may simply consider that the off-policy objective is itself a surrogate or that the gradients of the return are biased estimates based on past histories. Ideas introduced in this work also apply to other reinforcement learning techniques. Typically, for value-based RL with sparse-reward environments, convergence towards a value function outputting zero is expected with high probability. This is mostly due to the low probability of sampling non-zero rewards by Monte-Carlo. The discussions from Section 4 then apply, and a similar analysis can be performed.
@@
Our framework opens the door for further theoretical analysis, and the potential development of new criteria. We believe that deriving practical conditions on the exploration strategies, and the scheduling of the intrinsic return, for guaranteeing fast convergence should be the focus of attention. It could be achieved by bounding the policy improvement on expectation, which is nevertheless usually a hard task without strong assumptions. We furthermore believe that we provide a new lens on exploration necessary for interpreting and developing exploration strategies, in the sense of optimizing surrogate learning objective functions.
Acknowledgments
The authors thank Arnaud Delaunoy, Pascal Leroy, and Mathias Berger for valuable comments on this manuscript. Adrien Bolland and Gaspard Lambrechts gratefully acknowledge the financial support of the F.R.S.-FNRS.
Appendix A Reward Shaping and Exploration Strategies
As discussed in the manuscript, exploration strategies are reward-shaping strategies where the intrinsic reward bonuses are, among others, dependent on the policy parameters. This dependency makes the shaping strategies adaptive but makes the computation of gradients and the study of the learning objectives more complex. In this section, we study handcrafted reward-shaping strategies that have pseudoconcave and dense reward functions in the hill and maze environments. We then illustrate that the same criteria can be used to study these expert-knowledge based shaped rewards.
For the hill environment from Section 3, we illustrate in Figure 4(a) an intrinsic reward bonus making the sum of rewards in equation (2) concave. The corresponding learning objective has a unique maximum, which is part of the set $\Omega=\left\{\theta^{\prime}|\max_{\theta}J(\pi_{\theta})-J(\pi_{\theta^{\prime}})\leq\epsilon\right\}$ with $\epsilon=1$ and $\theta=(K,\sigma)$. It can be seen in Figure 4(b) where the global maximum in black is within the set $\Omega$ in green. Both, the $\epsilon$-coherence and the pseudoconcavity criteria are thus respected for $\epsilon=1$. Here, the intrinsic reward function is a simple function independent of the policy $\pi_{\theta}$. Finding such an intrinsic reward may be complex for other environments but the example underlines that exploration and reward shaping are mostly equivalent and that designing reward functions that are concave may help converging towards optimal policies.
(a) Alternative reward function.
(b) Learning objective.
Figure 4: In Figure 4(a), an alternative intrinsic reward function ensuring that the sum of rewards is a pseudoconcave function. In Figure 4(b), the contour function of the learning objective.
For the maze environment, the return $J(\pi_{\theta})$ is represented in black in Figure 5(a) together with the intrinsic return $J^{d}(\pi_{\theta})$ in green. The latter is the return of the dense handcrafted reward function $\rho^{d}(s,a)=(a-1)/2$ penalizing actions moving away from the target. In Figure 5(b), the corresponding learning objective function is shown. In the same experimental setting as in Section 4, we observe that the objective function is $\delta$-efficient for higher values of $\delta$ compared to the already-discussed learning objectives. Furthermore, the attraction criterion is respected for any value of $\delta$ as the unique global maxima of the learning objective, intrinsic return, and return are all equals.
(a) Return.
(b) Learning objectives.
(c) Improvement probabilities.
Figure 5: In Figure 5(a) the return of the maze environment is represented together with the intrinsic return of a dense handcrafted reward function. Figure 5(b) represents the corresponding learning objective and Figure 5(c) the probability that the REINFORCE estimates are positive.
Appendix B Minigrid Experiments
In this section, we introduce complex environments and parameterize policies with neural networks. In this context, it is impractical to naively compute and represent the objective functions and probability distributions for the different criteria. Therefore, we only evaluate the criteria along parameter trajectories, and extend the previous experimental setting.
We consider seven environments from the MiniGrid suite of environments (Chevalier-Boisvert et al., 2023), among others, designed for evaluating exploration strategies. In these environments, an agent moves in a maze and aims to reach a target position. To do so, the agent may choose actions that consist of turning left, turning right, moving forward, or staying idle. We consider two reward settings: the dense setting and the sparse setting. In the first, rewards of $-1$ are received for every non-idle move, and a reward of $1000$ is received upon reaching the target position. In the second setting, zero rewards are received everywhere, except upon reaching the target position, where a bonus of $1000$ is provided. In the dense setting, due to the action penalization incurred when moving, a policy outputting the idle action with probability one is locally optimal and has a return equal to zero. This is not (necessarily) the case in the second setting. We consider a discount factor of $\gamma=0.98$ and optimize a fully connected neural network taking as input the position pair and the orientation of the agent, and outputting a categorical distribution over actions. The network is composed of three hidden layers of $64$ neurons with ReLU activation functions.
In the dense reward setting, we optimize policies by maximizing three learning objective functions: $J(\pi_{\theta})$, $L^{a}(\theta)$, and $L^{s}(\theta)$, respectively with $\lambda_{a}=0.5$ and $\lambda_{s}=0.25$. For the last objective, the state-visitation density estimator is a ten-component Gaussian mixture model maximizing the likelihood of the sampled batch. The optimization is performed using the Adam update rule (Kingma & Ba, 2014), with REINFORCE ascent directions computed over 32 histories of constant length $T=100$, and with learning rate (step size) equal to $0.0005$. The length $T$ of the histories is chosen such that the realization value $T$ from a geometric distribution with success probability parameter $1-\gamma$ has at least a cumulative probability of $0.85$. In this setting, we illustrate the quasiconcavity criterion and the $\epsilon$-coherence criterion. In Figure 6, we provide the evolution of the return of the policies when optimizing the three objectives $J(\pi_{\theta})$, $L^{a}(\theta)$, and $L^{s}(\theta)$ for the different environments. For the MiniGrid-Empty-8x8-v0 and the MiniGrid-FourRooms-v0 environments, optimizing the return results in high-performance policies that do not stay idle. The other objectives also manage to find high-performing policies, but with a lower return. This phenomenon (assuming that the global optimum of each objective is found) illustrates the $\epsilon$-coherence criterion, where this $\epsilon$ value is the bound on the best policy that can be found when optimizing the learning objective. For the other environments, the policies resulting from the optimization of the return fall into local optima, namely ones where the policy chooses the idling action with probability one. When optimizing the learning objectives with exploration bonuses, the resulting policies no longer fall into the previous local optima. This result suggests that, along these parameter trajectories, the return $J(\pi_{\theta})$ has a local optimum (or saddle point), in opposition to the learning objective functions $L^{a}(\theta)$ and $L^{s}(\theta)$. The latter illustrates the validity of the pseudoconcavity criterion in that region of the parameter space. For the learning objective $L^{s}(\theta)$, the $\epsilon$-coherence criterion is respected for a small value of $\epsilon$, and the resulting policy manages to reach the target position. For the learning objective $L^{a}(\theta)$, the resulting policy does not reach the target position. We nevertheless hypothesize that it is not a real local optimum of the learning objective but a local optimum in the sense of equation (14).
Figure 6: Evolution of the return of policies during optimization in the dense minigrid environments. In blue, the return $J(\pi_{\theta})$ is optimized; in orange, the learning objective $L^{a}(\theta)$ is optimized; and in green, the learning objective $L^{s}(\theta)$ is optimized by performing Adam steps in REINFORCE directions. Note that the median, worst, and best cases over five runs are represented for the different curves. For the environments MiniGrid-Empty-8x8-v0 and MiniGrid-FourRooms-v0, optimizing each objective results in a policy that does not stay idle. These policies are initialized outside of the basin of attraction of the local optimum of the return. The coherence criteria can be observed as optimizing the learning objective with intrinsic exploration bonuses results in suboptimal policies. For the other environments, optimizing the return directly leads to policies that always choose to stay idle, and are thus locally optimal. Optimizing the learning objective with exploration allows us to escape from these local optima, illustrating the quasiconcavity criterion. It can be noted that when optimizing the objective $L^{a}(\theta)$, the $\epsilon$-coherence criterion is respected for a large value of $\epsilon$, making the resulting policy worse than when optimizing the return directly. On the contrary, the objective $L^{s}(\theta)$ appears to have $\epsilon$-coherence for a reasonable value of $\epsilon$.
In the previous experiments with the dense setting, the local optima exist due to the negative rewards associated to idle-actions. If we consider the sparse setting, we could then assume that directly optimizing the return is sufficient to find high-performing policies. It is not always the case and it can be justified using the efficiency and attraction criteria. We use the same parameters as in the previous set of experiments and we provide the evolution of the return of the policies when optimizing the three objectives $J(\pi_{\theta})$, $L^{a}(\theta)$, and $L^{s}(\theta)$, for the different environments in Figure 7. On the one hand, for most environments, whatever the learning objective, the resulting policy has a high return. Note that the $\epsilon$-coherence can again be illustrated where the policies resulting from the optimization of the return perform better than the others. On the other hand, for the MiniGrid-Empty-16x16-v0 and the MiniGrid-SimpleCrossingS11N5-v0 environments, a high performing policy can only be found when optimizing the learning objective $L^{s}(\theta)$. We illustrate that these results can be justified by the efficiency and attraction criteria in Figure 8. For each parameters obtained during the stochastic ascent steps on the return, we first estimate the probability of improving both objective functions by stochastic gradient ascent. These probabilities are used to compare learning objectives in terms of the efficiency criterion; the larger, the better. Then, in order to illustrate the attraction criterion, we estimate the probability of improving the return and the learning objective, both by gradient ascent steps on the learning objective, for each parameters obtained during the stochastic ascent steps on the objective $L^{a}(\theta)$ and $L^{s}(\theta)$. As far as $L^{a}(\theta)$ is concerned, all probabilities remain small as result that the optimization procedure converges fast towards a stationary point where the target goal is observed with negligible probability. The efficiency and attraction criteria are respected for negligible probabilities $\delta$, which is also a justification for the failure of converging towards good policies in the dense setting. For $L^{s}(\theta)$, on the contrary, and in both environments, the probability of improving the learning objective remains large for each parameter encountered when optimizing the return. The efficiency of the learning objective is much higher than that of the return in that part of the parameter space. Furthermore, the probability of improving the return when optimizing the learning objective, is small at the beginning and increases after some iterations. This indicates that once the policy has a sufficiently large intrinsic return, the attraction criterion is respected for a high value $\delta$.
@@ Appendix B Minigrid Experiments
Figure 7: Evolution of the return of policies during optimization in the sparse minigrid environments. In blue, the return $J(\pi_{\theta})$ is optimized; in orange, the learning objective $L^{a}(\theta)$ is optimized; and in green, the learning objective $L^{s}(\theta)$ is optimized performing Adam steps in REINFORCE directions. Note that the median, worst, and best cases over five runs are represented for the different curves. In most environments, a high-performing policy can be found by optimizing the return $J(\pi_{\theta})$. This results from its quasiconcavity. The $\epsilon$-coherence criterion can also be observed. However, for the MiniGrid-Empty-16x16-v0 and the MiniGrid-SimpleCrossingS11N5-v0 environments, a high-performing policy can only be found when optimizing the learning objective $L^{s}(\theta)$.
@@
Figure 8: For the MiniGrid-Empty-16x16-v0 and the MiniGrid-SimpleCrossingS11N5-v0 environments, and for both learning objective functions $L^{a}(\theta)$ and $L^{s}(\theta)$, we first represent the estimated probability of improving the return $J(\pi_{\theta})$ and the corresponding learning objective when following their REINFORCE gradient estimate. This value is estimated at each run of the optimization of the policy with learning objective $J(\pi_{\theta})$. Second, we represent the estimated probability of improving the corresponding learning objective and the return $J(\pi_{\theta})$ when following the REINFORCE gradient estimate of the learning objectives. These values are estimated at each run of the optimization of the policy with the learning objectives. The probabilities were estimated based on the frequencies of improving the objective functions by more than $0.2$ when following $5$ Adam ascent steps using REINFORCE update directions.

### v3

@@ 5 Conclusion
In conclusion, this research takes a step toward dispelling misunderstandings about exploration through the study of its effects on the performance of policy-gradient algorithms. More specifically, we distinguish two effects exploration has on the optimization. First, it modifies the learning objective to remove local extrema. Second, it modifies the gradient estimates and increases the likelihood that the update steps lead to improved returns. These two phenomena are studied through four criteria that we introduce and illustrate.
@@
These ideas apply to other direct policy optimization algorithms. Indeed, the four criteria do not assume any structure on the learning objective and can thus be straightforwardly applied to any objective function optimized by a direct policy search algorithm. In particular, for off-policy policy-gradient methods, we may simply consider that the off-policy objective is itself a surrogate or that the gradients of the return are biased estimates based on past histories. Ideas introduced in this work also apply to other reinforcement learning techniques. Typically, for value-based RL with sparse-reward environments, convergence to a value function that outputs zero is expected with high probability. This is mostly due to the low probability of sampling non-zero rewards via Monte Carlo. Similar discussion and analysis as in Section 4 then apply.
@@
Our framework opens the door to further theoretical analysis, and the potential development of new criteria. We believe that deriving practical conditions on the exploration strategies and the scheduling of the intrinsic return to guarantee fast convergence should be the focus of attention. It could be achieved by bounding the policy improvement in expectation, which is nevertheless usually a hard task without strong assumptions. We further believe that we provide a new lens on exploration that is necessary for interpreting and developing exploration strategies, in the sense of optimizing surrogate learning objective functions.
Appendix A Extended Experiments
In this section, we introduce complex environments and parameterize policies with neural networks. We then learn the policy parameters using policy gradients and intrinsic exploration bonuses. We discuss convergence results based on the four criteria introduced in the paper. In this context, it is impractical to naively evaluate the different criteria for each policy parameter, and we are restricted to a local evaluation of the criteria on the basis of statistics computed during optimization.
Let us adapt environments from the MiniGrid suite of environments (Chevalier-Boisvert et al.,, 2023). Each environment corresponds to a maze in which an agent moves, aiming to reach a target position. The agent may choose as actions to turn left, turn right, move forward, or stay idle. We consider two reward settings: dense and sparse. In the first, rewards of 1-1 are received for every non-idle move, and a reward of $1000$ is received upon reaching the target position. In the second setting, zero rewards are received everywhere, except upon reaching the target position, where a bonus of $1000$ is provided. We consider $\gamma=0.98$.
In the following, the policy is parameterized with a fully connected neural network taking as input the position pair and the orientation of the agent, and outputting a categorical distribution over actions. The network is composed of three hidden layers of $64$ neurons with ReLU activation functions. The policy is optimized in the different environments and reward settings. To that end, we consider three learning-objective functions: $J(\pi_{\theta})$, $L^{a}(\theta)$, and $L^{s}(\theta)$, respectively with $\lambda_{a}=0.5$ and $\lambda_{s}=0.25$. For the last objective, the state-visitation density estimator is a ten-component Gaussian mixture model that maximizes the likelihood of the sampled batch. The optimization is performed using the Adam update rule (Kingma and Ba,, 2014), with REINFORCE ascent directions computed over $32$ histories of constant length $T=100$, and with a learning rate (step size) equal to $0.0005$. The length $T$ of the histories is chosen such that the realized value of $T$ from a geometric distribution with success probability parameter $1-\gamma$ has at least a cumulative probability of $0.85$.
We first consider the environments in the dense-reward setting and discuss convergence in light of the pseudoconcavity and $\epsilon$-coherence criteria. In Figure 4, we provide the evolution of the return of the policies when optimizing the three objectives $J(\pi_{\theta})$, $L^{a}(\theta)$, and $L^{s}(\theta)$ for the different environments. On the one hand, in the MiniGrid-Empty-8x8-v0 and MiniGrid-FourRooms-v0 environments, optimizing the return provides high-performing policies, and the policies resulting from the optimization of the other learning objectives are suboptimal. The difference in performance is justified by the $\epsilon$-coherence criterion, where $\epsilon$ is the bound on the best policy that can be found when optimizing the learning objective (assuming that the global optimum of each objective is found). On the other hand, in the other environments, optimizing the return often leads to convergence to locally optimal parameters (or saddle points). These parameters correspond to idling policies and are locally optimal due to the action penalization incurred when moving. The policies obtained when optimizing $L^{a}(\theta)$ and $L^{s}(\theta)$ no longer fall into the previous local optima. The latter illustrates the validity of the pseudoconcavity criterion in that region of the parameter space. Note that when optimizing $L^{a}(\theta)$, the policy tends to converge towards another local optimum where actions are uniformly distributed. We nevertheless hypothesize it is local in the sense of equation (14).
@@ Appendix A Extended Experiments
Figure 4: Return of policies during optimization in the dense minigrid environments. In blue, the return $J(\pi_{\theta})$ is optimized; in orange, the learning objective $L^{a}(\theta)$ is optimized; and in green, the learning objective $L^{s}(\theta)$ is optimized by performing Adam steps in REINFORCE directions. Note that the median, worst, and best cases over five runs are represented for the different curves.
In the previous experiments, the local optima arise due to the negative rewards associated with idle actions. If we consider the sparse-reward setting, we could then assume that directly optimizing the return is sufficient to find high-performing policies. We use the same parameters as previously and provide the evolution of the return of the policies when optimizing the three objectives $J(\pi_{\theta})$, $L^{a}(\theta)$, and $L^{s}(\theta)$ for the different environments in Figure 5. On the one hand, for most environments, regardless of the learning objective, the resulting policy has a high return. Note that the $\epsilon$-coherence can again be illustrated, as the policies resulting from the optimization of the return perform better than the others. On the other hand, for the MiniGrid-Empty-16x16-v0 and MiniGrid-SimpleCrossingS11N5-v0 environments, a high-performing policy can only be found when optimizing the learning objective $L^{s}(\theta)$. We illustrate in the following that these results can be justified by the efficiency and attraction criteria.
Figure 5: Return of policies during optimization in the sparse minigrid environments. In blue, the return $J(\pi_{\theta})$ is optimized; in orange, the learning objective $L^{a}(\theta)$ is optimized; and in green, the learning objective $L^{s}(\theta)$ is optimized performing Adam steps in REINFORCE directions. Note that the median, worst, and best cases over five runs are represented for the different curves.
In Figure 6(a), we show the probability of improving the return $J(\pi_{\theta})$ and the learning objective $L^{s}(\theta)$ for each parameter obtained during their respective optimization by stochastic gradient ascent steps (i.e., SGA following respectively $\hat{g}\approx\nabla_{\theta}J(\pi_{\theta})$ and $\hat{d}\approx\nabla_{\theta}L^{s}(\theta)$). The improvement probability is higher for the learning objective than for the return. In other words, for these parameters, the efficiency $\delta$ of the learning objective is much higher than that of the return, which justifies the failure of optimization without exploration in Figure 5. In Figure 6(b), we again show the probability of improving the return $J(\pi_{\theta})$ and the learning objective $L^{s}(\theta)$, but for each parameter obtained during the stochastic gradient ascent steps on $L^{s}(\theta)$ (i.e., SGA following $\hat{d}\approx\nabla_{\theta}L^{s}(\theta)$). The probability of improving the return when optimizing the learning objective is small at the beginning and increases after some iterations. This indicates that after some iterations, the attraction criterion holds for a large $\delta$. The return thus eventually starts increasing, and optimizing $L^{s}(\theta)$ provides a high-performing policy, as can be seen in Figure 5.
(a) SGA following $\hat{g}$ and $\hat{d}$
(b) SGA following $\hat{d}$
@@
Figure 6: In Figure 6(a), we show the estimated probability of improving the return $J(\pi_{\theta})$ and the learning objective $L^{s}(\theta)$ when following their REINFORCE gradient estimates. In Figure 6(b), we show the estimated probability of improving the return $J(\pi_{\theta})$ and the learning objective $L^{s}(\theta)$ when following the REINFORCE gradient estimate of the learning objective $L^{s}(\theta)$. Probabilities were estimated based on the frequencies of improving the objective functions by more than $0.2$ after following 55 Adam ascent steps.